### v1

@@ 1 Introduction
where $P_{k}$ is the space of polynomials $p$ of degree less or equal than $n$ and unitary lowest order coefficient, $\Lambda(B)$ is the spectrum of $B$. The convergence behaviour is not fully characterized by the condition number of $B$, but by its spectral distribution, and that an adaptive bound as in [19] would be more descriptive of the convergence behaviour of the method. However, bound (4) indicates that if the condition number of the matrix is very close to 1, then the method will converge very fast independently of the spectral distribution of $B$, and this is enough for our discussion. The GMRES method is very different from this point of view. Indeed, its convergence is determined by the minimal polynomial of the matrix ${A}$ [26], and the spectral properties of $A$ are not enough to fully characterize the convergence of the method. In fact, it can be proven that any nonincreasing GMRES convergence curve is possible for any eigenvalue distribution, [16, 6]. This implies that preconditioning strategies for this method are often derived by heuristic processes. It is worth mentioning that if the matrix $A$ is nearly orthogonally diagonalizable the convergence of GMRES it is well understood in terms of the spectrum of $A$, [26, Theorem 35.2].
@@ 5 Conclusions
    Mixed reformulations of (20) can be used to escape the fact that $A^{T}TA$ is a dense matrix. Specifically, we aim to study this problem from the point of view of its Schur complement [29]. Such reformulation has many similarities to the discontinuous Petrov–Galerkin method [4, 7], which we plan to highlight in future work.
@@ Acknowledgements
The authors would like to thank P. Brubeck, P. E. Farrell, A. J. Wathen, C. Parker, and S. Zampini for the fruitful discussions and invaluable suggestions, while writing this paper.

### v2

@@ 1 Introduction
where $P_{k}$ is the space of polynomials $p$ of degree less or equal than $n$ and unitary lowest order coefficient, $\Lambda(B)$ is the spectrum of $B$. The convergence behaviour is not fully characterized by the condition number of $B$, but by its spectral distribution, and that an adaptive bound as in [20] would be more descriptive of the convergence behaviour of the method. However, bound (4) indicates that if the condition number of the matrix is very close to 1, then the method will converge very fast independently of the spectral distribution of $B$, and this is enough for our discussion. The GMRES method is very different from this point of view. Indeed, its convergence is determined by the minimal polynomial of the matrix ${A}$ [27], and the spectral properties of $A$ are not enough to fully characterize the convergence of the method. In fact, it can be proven that any nonincreasing GMRES convergence curve is possible for any eigenvalue distribution, [17, 7]. This implies that preconditioning strategies for this method are often derived by heuristic processes. It is worth mentioning that if the matrix $A$ is nearly orthogonally diagonalizable the convergence of GMRES it is well understood in terms of the spectrum of $A$, [27, Theorem 35.2].
@@ 4 Functional Analytic Perspective
Remark 4.9 (Relations with LSQ-FEM).
There is an intimate relation between (20) and the Least SQuares Finite Element Method (LSQ-FEM) [4]. In fact, (20) can be recast as a minimization problem, i.e
(47) $\mathbf{x}=\underset{\mathbf{w}\in V_{h}}{\text{argmin}}\;J(\mathbf{w};\mathbf{b}),\qquad J(\mathbf{w};\mathbf{b})\coloneqq\left\lVert A\mathbf{w}-\mathbf{b}\right\rVert_{T}^{2}.$
This effectively reduces the discretisation assosciated with the normal equation (20) to a rather non-standard LSQ-FEM discretisation. We stress that this is non-standard because it is typical in LSQ-FEM to view the least square functional as a residual associated with the strong form of the PDE or with one of its first order reformulation. Yet a key distinction between the strategy discussed here and a non-standard least square finite elements formulation lies in the fact that while a LSQ-FEM method construct a variational formulation by selecting appropriate norms, function spaces and residual, the strategy here seeks to identify the continuous PDE discretised by the operator $A^{T}TA$ for different choices of inner product $T$ and pick the inner product $T$ resulting in the most convenient PDE to precondition.
All codes used to generate the results presented in this section can be found in the repository https://github.com/UZerbinati/normalPreconditioning.
@@ 5 Conclusions
    Mixed reformulations of (20) can be used to escape the fact that $A^{T}TA$ is a dense matrix. Specifically, we aim to study this problem from the point of view of its Schur complement [30]. Such reformulation has many similarities to the discontinuous Petrov–Galerkin and least squares finite element methods [8, 5, 4], which we plan to highlight in future work.
@@ Acknowledgements
The authors would like to thank P. Brubeck, P. E. Farrell, J. Málek, C. Parker, V. Simoncini, A. J. Wathen, and S. Zampini for the fruitful discussions and invaluable suggestions, while writing this paper.